OpenStudy (anonymous):

of all rectangles that have a perimeter of 34 ft, find the dimension with the largest area? What is its area?

OpenStudy (freckles):

calculus or algebra?

OpenStudy (freckles):

Because we could go either route here.

OpenStudy (anonymous):

calculus

OpenStudy (freckles):

First all the perimeter is given to you. We have 2L+2W=34 We want to maximize the area, A=LW

OpenStudy (freckles):

So solve 2L+2W=34 for either L or W. Doesn't matter which one you choose.

OpenStudy (anonymous):

how about L

OpenStudy (freckles):

Your current objective is to write A in terms of one variable (be it L or W but not both)

OpenStudy (freckles):

So what do you get when you solve 2L+2W=34

OpenStudy (freckles):

for L

OpenStudy (anonymous):

L=17-W?

OpenStudy (freckles):

So recall A=LW so we now have A=(17-W)W Distribute there then find A'

OpenStudy (freckles):

A' being dA/dW

OpenStudy (anonymous):

why is there a W on the outside

OpenStudy (anonymous):

equation

OpenStudy (freckles):

A=L*W

OpenStudy (freckles):

L=(17-W)

OpenStudy (freckles):

so A=(17-W)*W

OpenStudy (freckles):

I replaced L in A=LW with (17-W)

OpenStudy (anonymous):

A=17-2W ?

OpenStudy (freckles):

That is A' or A?

OpenStudy (freckles):

I think that you mean A'=17-2W

OpenStudy (anonymous):

A'

OpenStudy (freckles):

Now find your critical numbers by finding when A'=0

OpenStudy (freckles):

17-2W=0 when W=?

OpenStudy (anonymous):

W=8.5

OpenStudy (freckles):

Now you can find L These will be the dimensions that maximize the area and therefore you can also find the maximum area you can verify this is a max by finding A'' A''=-2<0 therefore W=8.5 is a max

OpenStudy (freckles):

well W=8.5 is where the max occurs *

OpenStudy (anonymous):

So am i suppose to take plug in the W into the original equation or take the second derivative of the equation A'=17-2W?

OpenStudy (anonymous):

so i plugged 8.5 back into the original equation and L= 8.5 as well. That doesn't really make sense because it's a rectangle not a square

OpenStudy (freckles):

You need find L for when W=8.5 remember the relationship between W and L was L=17-W

OpenStudy (freckles):

Right L=8.5

OpenStudy (freckles):

rectangles are squares when their side lengths are the same

OpenStudy (freckles):

And a square is what gives a max area

OpenStudy (freckles):

So the actual max area is 8.5*8.5

OpenStudy (anonymous):

I'm still not following. How is a rectangle same as a square?

OpenStudy (freckles):

a rectangle is a square when the lengths of a rectangle are the same

OpenStudy (freckles):

therefore a rectangle can be a square

OpenStudy (freckles):

but not all rectangles are squares

OpenStudy (freckles):

so some rectangles are squares

OpenStudy (anonymous):

okay so the area is 72.25

OpenStudy (freckles):

anymore questions?

OpenStudy (freckles):

I can also show you algebraically you will get the same dimensions for your rectangle.

OpenStudy (anonymous):

Yes i have another question

OpenStudy (freckles):

\[2L+2W=34 \\ L+W=17 \\ A=LW=(17-W)W \\A=-W^2+17W \\ \text{ A is a parabola open downward, therefore it has a max and not a min}\\ A=-(W^2-17W) \\A=-(W^2-17W+(\frac{17}{2})^2)+(\frac{17}{2})^2 \\ A=-(W-\frac{17}{2})^2+(\frac{17}{2})^2 \\ \text{ the vertex where the max occurs is } (\frac{17}{2}, (\frac{17}{2})^2) \text{ So we have a \max area when } \\ W=\frac{ 17}{2} \text{ and } A=(\frac{17}{2})^2\] And so L=17/2

OpenStudy (freckles):

so all squares are rectangles but not all rectangles are squares

OpenStudy (freckles):

what is the other question

OpenStudy (anonymous):

From a thin piece of cardboard 20 in by 20 in, square corners are cut out so that the sides can be folded up to make a box. What dimensions will yield a box of maximum volume? What is the maximum volume?

OpenStudy (freckles):

Ok a drawing would help here

OpenStudy (freckles):

Can you draw it?

OpenStudy (anonymous):

I'm sorry but there is no picture provided in my math textbook :(

OpenStudy (freckles):

Act like you are constructing an open box from some cardboard

OpenStudy (freckles):

And construct it exactly like it says

OpenStudy (freckles):

No lol I was asking you to draw the picture

OpenStudy (freckles):

I can draw the picture The construction of the picture is for you

OpenStudy (anonymous):